# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

### v2

@@ 4.3 Attack Stealthiness
We visualize some poisoned samples generated by different backdoor attack methods in Fig. 3. As observed, PointBA-I introduces noticeable anomalies compared to clean samples, significantly increasing its detectability. Although PointBA-O and IRBA generate natural-looking poisoned samples by applying transformation, they still cause considerable deformations in shape. Compared to IBAPC, SPBA selectively perturbs the flower region while preserving the smooth structure of the vase, making the modifications less perceptible to human observers. Additional visualization results are provided in the Supplementary. Furthermore, we conduct a human perceptual study to evaluate the stealthiness of the attack. Following the setting in IRBA [4], a mixture of clean and poisoned samples is presented to 20 participants. Each participant rates the probability of a sample being maliciously modified on a 5-point scale, where lower scores indicate higher stealthiness. Tab. 3 shows that SPBA achieves the best stealthiness with the lowest score.
@@
4.4 Attack Efficiency
As shown in the Eq. (6), the computational cost of spectral attacks is highly dependent on the trigger size. And we further report the computational cost of different spectral attacks (with a batch size of 64 during the inference stage) in Tab. 4. The prior spectral attack IBAPC uses a large trigger of size $\xi^{\text{IBAPC}}\in\mathbb{R}^{1024\times 3}$, resulting in a substantial computational cost of 939.72 MFLOPs, which severely hinders its efficiency and practicality for deployment. In contrast, although our proposed SPBA perturbs an average of 452 points, which account for 44% of the sample, our patch-based spectral trigger is much smaller ($\xi^{SPBA}\in\mathbb{R}^{32\times 3}$). The computational cost of SPBA is only 14.74 MFLOPs, which is comparable to that of a single fully connected layer. Furthermore, similar to IBAPC, we construct a sample-based spectral trigger using the FPS algorithm (denoted as FPSP, $\in\mathbb{R}^{452\times 3}$) that perturbs the same number of points as SPBA. However, it significantly increases the computational cost by a factor of 11.54 compared to our patch-based trigger. Moreover, it results in weaker attack performance (shown in Tab. 7), highlighting the advantage of our patch-wise design. These results demonstrate the efficiency and effectiveness of our proposed patch-wise spectral trigger.
Table 5: Resistance of backdoor attacks to varied data augmentations.
@@ 4.4 Attack Efficiency
4.5 Resistance to Defense Methods
Resistance to Data Augmentation. Since data augmentation is widely employed to enhance the resistance of classification models, we introduce six types of augmentations to evaluate their impact on backdoor attack performance: 1) Rotation: random rotation around the z-axis with a maximum angle of $10^{\circ}$; 2) Rotation3D: random rotation up to $10^{\circ}$ along all three axes; 3) Scaling: scaling within a range of 0.5 to 1.5; 4) Shift: shifting randomly between -0.1 and 0.1 along each axis; 5) Dropout: removing up to 20% of points; and 6) Jitter: introducing noise sampled from $N(0,0.02)$. The experiments are conducted on the ModelNet40 dataset using the PointNet++ model. As shown in Tab. 5, different augmentations affect attack methods to varying degrees. Due to the high visibility of the trigger of PointBA-I, augmentations have minimal impact, as the trigger remains distinct. However, for rotation-based PointBA-O, applying Rotation significantly reduces attack effectiveness. For IRBA, which relies on local transformations, augmentations such as Rotation3D, Dropout and Jitter cause a notable decline in attack performance, with ASR dropping by up to 8.23%. Conversely, for optimization-based attacks like IBAPC and our proposed SPBA, only Jitter leads to a slight ASR reduction, while other augmentations have negligible effects, demonstrating their resistance against most augmentations.
Resistance to SOR. Statistical Outlier Removal (SOR) [36] is designed to eliminate outliers based on the local neighborhood density using the KNN algorithm. In our implementation, we set KNN to use 40 nearest neighbors and set the standard deviation threshold to 1. We apply it during inference on PointNet++ models trained on ModelNet40. The results show that only PointBA-I experiences a significant drop in ASR, decreasing from 98.87% to 75.44%, while the ASR reduction for other attacks remains within 3%. This suggests that SOR serves as an effective defense specifically against injection-based attacks while having a limited impact on other attack methods.
Resistance to PointCRT. PointCRT [9] utilizes a nonlinear classifier to detect the abnormal corruption resistance of poisoned samples. The F1 score (F1) and Area Under the Curve (AUC) are employed to evaluate the ability of PointCRT to detect the attack, where lower values indicate that the attack is more resistant to detection by PointCRT. As shown in Tab. 6, SPBA outperforms almost all the attacks except PointBA-O, which exhibits limited attack performance and is vulnerable to data augmentations.
Table 6: Detect accuracy of PoinCRT on varies attacks.
@@ 4.5 Resistance to Defense Methods
Table 8: Effect of selected patch number $m$.
@@
4.6 Ablation Study
Effect of Patch Selection Strategies. In our default setting, we select the 16 patches with the highest patch imperceptibility scores (PIS), denoted as H-PSI. To evaluate the effectiveness of this selection strategy, we compare H-PSI with three alternative patch selection strategies: (1) L-PSI: Select the patches with the lowest PIS; (2) Random: Randomly select the patches; (3) FPSP: Use the FPS algorithm to sample the same number of points as H-PSI. We evaluate their average performance on ModelNet40. As shown in Tab. 7, the H-PSI strategy achieves the highest attack effectiveness and the strongest stealthiness. This result underscores the advantage of selecting high-imperceptibility patches for local spectral trigger injection.
Effect of Selected Patch Number $m$. In this section, we experiment with injecting the spectral trigger into varying numbers $m$ of selected patches during the training process. The experiments are conducted on ModelNet40 using PointNet++. As shown in Tab. 8, as $m$ increases from 1 to 16, the CD value gradually decreases, indicating improved stealthiness. This suggests that selecting more patches within a reasonable range helps distribute perturbations more naturally, enhancing imperceptibility. However, when $m$ further increases to 32, where almost all points are perturbed, CD rises to 0.12, indicating that excessive modifications start to introduce noticeable distortions, reducing stealthiness. Based on these observations, we set $m=16$ as the default configuration, as it achieves the optimal balance between imperceptibility (lowest CD), strong attack effectiveness, and benign accuracy.
@@ 5 Conclusion
Appendix
Appendix A Deformation by Spectral Perturbation
As illustrated in Figure 4, low-spectrum perturbations cause significant deformation to the patch of empennage, while high-spectrum perturbations introduce excessive outliers. Perturbations applied within intermediate spectral bands mitigate outliers while preserving the overall shape, presenting a key challenge for the optimal design of spectral triggers. To address this, we propose an optimizable spectral trigger $\xi$ that strikes a balance between attack effectiveness and stealthiness.
Figure 4: Comparison of original and perturbed patches. The second row shows the spectral features of the original patch and spectral perturbations across different spectrum bands.
Appendix B Resistance to the Gradient-based Method
As many studies on backdoor attacks assess trigger stealthiness using gradient-based methods [22, 17], we employ a point cloud saliency map [35] to identify salient points that contribute most to the model’s decision-making. Specifically, 40 points with the highest saliency scores are marked in red. As illustrated in Figure 5, the spherical trigger of PointBA-I is easily detected. In contrast, our proposed SPBA exhibits a saliency distribution similar to the clean model, indicating enhanced stealthiness and reduced detectability.
Appendix C Ablation Study
C.1 Effect of the Target Class
In this section, we investigate the effect of different target classes on our attack performance. Specifically, we select the top 8 categories from ModelNet40 as target classes and conduct experiments using PointNet++. As shown in Figure 7, the choice of target class has a minimal effect on both Attack Success Rate (ASR) and Benign Accuracy (BA), with fluctuations not exceeding 1%. These results demonstrate the robustness of our approach across various target label settings, ensuring its consistency in different attack scenarios.
Figure 6: Visualization on more poisoned samples.
Figure 7: Effect of the target class selection.
C.2 Effect of the Poisoning Rate
We explore the impact of the poisoning rate on our attack by conducting experiments on ModelNet40 using PointNet++. As shown in Figure 8, ASR consistently increases as the poisoning ratio rises, peaking at 98.42%. BA exhibits minor fluctuations but remains relatively stable as the poisoning rate increases. Notably, even at a low poisoning rate of 2.5%, our approach achieves an ASR of 96.27%, underscoring its effectiveness even with limited poisoned data.
Figure 8: The effect of the poisoning rate.
Appendix D Visualization on More Poisoned Samples
In Figure 6, we provide additional visualization results that highlight the exceptional stealthiness of our attack method. Our approach not only maintains the overall structural integrity of the poisoned samples but also ensures that smooth regions, such as the bottle body, chair legs, and car body, remain undisturbed. This careful preservation further enhances the imperceptibility of the attack.

### v3

@@ 5.3 Attack Stealthiness
We visualize some poisoned samples generated by different backdoor attack methods in Fig. 4. As observed, PointBA-I introduces noticeable anomalies compared to clean samples, significantly increasing its detectability. Although PointBA-O and IRBA generate natural-looking poisoned samples by applying transformation, they still cause considerable deformations in shape. Compared to IBAPC, SPBA selectively perturbs the flower region while preserving the smooth structure of the vase, making the modifications less perceptible to human observers.
Furthermore, we conduct a human perceptual study to assess the stealthiness of the proposed attack. Following the protocol in IRBA [7], we present a mixture of clean and poisoned samples to 20 participants, all of whom are over 18 years old, major in computer science, and have prior knowledge of machine learning and computer vision. Each participant is shown 10 groups of multi-view renderings of point clouds from distinct object classes. Each group consists of one clean reference sample together with five attack types of its poisoned version arranged in random order. Participants are instructed to assign a score $\in\{1,2,3,4,5\}$ to indicate the likelihood that the poisoned samples in the group have been modified. A lower score reflects a lower perceived likelihood, suggesting that the poisoned samples are less perceptible. The average score for each attack method is reported in Table 3. The results demonstrate that SPBA achieves the highest level of stealthiness, yielding the lowest perceptibility score among all compared methods.
Figure 4: Visual comparison of poisoned samples from different backdoor attack methods. Our proposed SPBA preserves structural integrity and ensures smooth regions (e.g., the vase or car roof) remain undisturbed, enhancing imperceptibility.
@@
5.4 Attack Efficiency
As shown in the Eq. (6), the computational cost of spectral attacks is highly dependent on the trigger size. And we further report the computational cost of different spectral attacks (with a batch size of 64 during the inference stage) in Table 4. The prior spectral attack IBAPC uses a large trigger of size $\xi^{\text{IBAPC}}\in\mathbb{R}^{1024\times 3}$, resulting in a substantial computational cost of 939.72 MFLOPs, which severely hinders its efficiency and practicality for deployment. In contrast, although our proposed SPBA perturbs an average of 452 points, which account for 44% of the sample, our patch-based spectral trigger is much smaller ($\xi^{SPBA}\in\mathbb{R}^{32\times 3}$). The computational cost of SPBA is only 14.74 MFLOPs, which is comparable to that of a single fully connected layer. Furthermore, similar to IBAPC, we construct a sample-based spectral trigger using the FPS algorithm (denoted as FPSP, $\in\mathbb{R}^{452\times 3}$) that perturbs the same number of points as SPBA. However, it significantly increases the computational cost by a factor of 11.54 compared to our patch-based trigger. Moreover, it results in weaker attack performance (shown in Table 8), highlighting the advantage of our patch-wise design. These results demonstrate the efficiency and effectiveness of our proposed patch-wise spectral trigger.
@@ 5.4 Attack Efficiency
5.5 Resistance to Defense Methods
Resistance to Data Augmentation. Since data augmentation is widely employed to enhance the resistance of classification models, we introduce six types of augmentations to evaluate their impact on backdoor attack performance: 1) Rotation: random rotation around the z-axis with a maximum angle of $10^{\circ}$; 2) Rotation3D: random rotation up to $10^{\circ}$ along all three axes; 3) Scaling: scaling within a range of 0.5 to 1.5; 4) Shift: shifting randomly between -0.1 and 0.1 along each axis; 5) Dropout: removing up to 20% of points; and 6) Jitter: introducing noise sampled from $N(0,0.02)$. The experiments are conducted on the ModelNet40 dataset using the PointNet++ model. As shown in Table 5, different augmentations affect attack methods to varying degrees. Due to the high visibility of the trigger of PointBA-I, augmentations have minimal impact, as the trigger remains distinct. However, for rotation-based PointBA-O, applying Rotation significantly reduces attack effectiveness. For IRBA, which relies on local transformations, augmentations such as Rotation3D, Dropout and Jitter cause a notable decline in attack performance, with ASR dropping by up to 8.23%. Conversely, for optimization-based attacks like IBAPC and our proposed SPBA, only Jitter leads to a slight ASR reduction, while other augmentations have negligible effects, demonstrating their resistance against most augmentations.
Table 5: Resistance of backdoor attacks to varied data augmentations and SOR.
@@ 5.5 Resistance to Defense Methods
SOR 90.36 46.47 89.55 95.46 90.40 97.30 89.67 95.13 89.42 97.36
Resistance to SOR. Statistical Outlier Removal (SOR) [48] is designed to eliminate outliers based on the local neighborhood density using the KNN algorithm. In our implementation, we set KNN to use 60 nearest neighbors and set the standard deviation threshold to 0.5. We apply it during inference on PointNet++ models trained on ModelNet40. The results in Table 5 show that only PointBA-I experiences a significant drop in ASR, decreasing from 98.87% to 36.50%, while the ASR reduction for other attacks remains within 2%. This suggests that SOR serves as an effective defense specifically against injection-based attacks while having a limited impact on other attack methods.
Resistance to the Gradient-based Method. As many studies on backdoor attacks assess trigger stealthiness using gradient-based methods [30, 24], we employ a point cloud saliency map [47] to identify salient points that contribute most to the model’s decision-making. Specifically, 40 points with the highest saliency scores are marked in red. As illustrated in Fig. 5, the spherical trigger of PointBA-I is easily detected. In contrast, our proposed SPBA exhibits a saliency distribution similar to the clean model, indicating enhanced stealthiness and reduced detectability.
Resistance to PointCRT. PointCRT [15] utilizes a nonlinear classifier to detect the abnormal corruption resistance of poisoned samples. The F1 score (F1) and Area Under the Curve (AUC) are employed to evaluate the ability of PointCRT to detect the attack, where lower values indicate that the attack is more resistant to detection by PointCRT. As shown in Table 6, SPBA outperforms almost all the attacks except PointBA-O, which exhibits limited attack performance and is vulnerable to data augmentations.
Table 6: Performance of PoinCRT on varies attacks.
@@
Resistance to FLARE. FLARE [13] aggregates abnormal activations across hidden layers and adaptively selects the most stable subspace to divide samples into two clusters, detecting the more stable cluster as poisoned samples. Since this method is independent of input modality, it could also be applied to point cloud backdoor sample detection. Specifically, we evaluated the method on backdoored DGCNN models trained on ModelNet40. The true positive rate (TPR) reflects the proportion of correctly identified poisoned training samples, with a lower TPR indicating weaker detection capability. Experimental results in Table 7 show that our approach consistently yields the lowest TPR, demonstrating strong resistance against FLARE.
Table 7: Detect accuracy of FLARE on varies attacks.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
TPR $\downarrow$ 56.93 62.01 47.97 38.01 29.78
5.6 Ablation Study
Effect of Patch Selection Strategies. In our default setting, we select the 16 patches with the highest patch imperceptibility scores (PIS), denoted as H-PSI. To evaluate the effectiveness of this selection strategy, we compare H-PSI with three alternative patch selection strategies: (1) L-PSI: Select the patches with the lowest PIS; (2) Random: Randomly select the patches; (3) FPSP: Use the FPS algorithm to sample the same number of points as H-PSI. We evaluate their average performance on ModelNet40. As shown in Table 8, the H-PSI strategy achieves the highest attack effectiveness and the strongest stealthiness. This result underscores the advantage of selecting high-imperceptibility patches for local spectral trigger injection.
@@ 5.6 Ablation Study
Effect of Selected Patch Number $m$. In this section, we experiment with injecting the spectral trigger into varying numbers $m$ of selected patches during the training process. The experiments are conducted on ModelNet40 using PointNet++. As shown in Table 9, as $m$ increases from 1 to 16, the CD value gradually decreases, indicating improved stealthiness. This suggests that selecting more patches within a reasonable range helps distribute perturbations more naturally, enhancing imperceptibility. However, when $m$ further increases to 32, where almost all points are perturbed, CD rises to 0.12, indicating that excessive modifications start to introduce noticeable distortions, reducing stealthiness. Based on these observations, we set $m=16$ as the default configuration, as it achieves the optimal balance between imperceptibility (lowest CD), strong attack effectiveness, and benign accuracy.
Table 9: The effect of selected patch number $m$.
@@
Effect of the Patch Size. We conduct additional experiments on ModelNet40 using PointNet++ to further investigate the effect of patch size. To ensure that the total number of perturbed points remains approximately constant, when the patch size is doubled, the number of selected patches is reduced to half, and vice versa. From Table 10, ASR attains its maximum at $k_{g}=32$ and CD is minimized at this setting, offering the best balance of attack success and stealth. The benign accuracy is still high (89.42%), ranking second among the evaluated patch sizes.
Table 10: The effect of patch size $k_{g}$
$k_{g}$ 16 32 64 128
BA 89.30 89.42 89.75 89.26
ASR 97.00 97.97 97.64 97.73
CD 0.13 0.11 0.18 0.15
Effect of the Target Class. In this section, we investigate the effect of different target classes on our attack performance. Specifically, we select the top 8 categories from ModelNet40 as target classes and conduct experiments using PointNet++. As shown in Fig. 6(a), the choice of target class has a minimal effect on both Attack Success Rate (ASR) and Benign Accuracy (BA), with fluctuations not exceeding 1%. These results demonstrate the robustness of our approach across various target label settings, ensuring its consistency in different attack scenarios.
(a) The effect of the target class selection.
(b) The effect of poisoning rate.
Figure 6: The effect of the target class selection and poisoning rate.
Effect of the Poisoning Rate. We explore the impact of the poisoning rate on our attack by conducting experiments on ModelNet40 using PointNet++. As shown in Fig. 6(b), ASR consistently increases as the poisoning ratio rises. BA exhibits minor fluctuations but remains relatively stable as the poisoning rate increases. Notably, even at a low poisoning rate of 2.5%, our approach achieves an ASR of 96.27%, underscoring its effectiveness even with limited poisoned data.
Effect of the Regulation Loss Weights. In this section, we investigate the effect of the regulation loss weights in Eq. 10, where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ control the contributions of $L_{ED}$, $L_{CD}$, and $L_{HD}$, respectively. As shown in Table 11, increasing $\lambda_{1}$ gradually enhances the stealthiness (reflected by lower CD values), but at the cost of a decreasing ASR. Notably, when $\lambda_{1}=1$, the attack achieves a relatively good trade-off between aggressiveness and stealthiness. Moreover, our proposed SPBA algorithm demonstrates robustness with respect to the hyperparameters $\lambda_{2}$ and $\lambda_{3}$. Based on these observations, we ultimately set $\lambda_{2}=5$ and $\lambda_{3}=1$ to achieve a favorable balance between attack effectiveness and stealthiness.
Table 11: The effect of the regulation loss weights
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ BA ASR CD
1 5 1 89.42 97.97 0.11
0.1 5 1 89.72 99.39 0.72
0.5 5 1 89.47 98.50 0.19
2 5 1 87.64 92.70 0.08
1 1 1 89.95 97.17 0.15
1 3 1 89.34 97.77 0.11
1 5 0.1 89.74 96.96 0.12
1 5 0.5 89.62 97.44 0.11
1 5 2 89.52 97.71 0.11